# Trojan Cleansing with Neural Collapse

### v2

@@ 4.1 Methodology
This tension motivates an algorithm for cleansing a network of any potential trojan triggers, which we name ETF-FT. ETF-FT aims to take a fully trained, possibly trojaned neural network and remove trojan triggers, if any exist, while maintaining the network’s good performance on clean test data. The effects of a trojan attack are most apparent in Fig. 6(a), which shows that the weights of the final layer of a benign neural network exhibit a much greater degree of symmetry than those of an analogous trojaned network. So, as specified in Algorithm 1, we proceed by first over-writing the weights of the final fully-connected layer to a randomly generated simplex ETF, freezing those weights, and then fine-tuning the remaining model parameters on a small subset of clean data. Notably, this process works robustly with limited data, and the fine-tuning data does not necessarily need to come from the original training set.
PreAct-ResNet18[13] ViT-B/16[1]
5% Clean data 1% Clean data 5% 1%
Method ACC ASR ACC ASR ACC ASR ACC ASR
Original 88.95 95.06 88.95 95.06 95.43 87.06 95.43 87.06
FT $91.63\pm 0.18$ $10.03\pm 7.23$ $90.20\pm 0.22$ $22.04\pm 25.38$ 96.06 20.46 95.74 36.87
NCleanse[29] $91.57\pm 0.1$ $\textbf{0.98}\pm 0.12$ $89.90\pm 0.34$ $\textbf{1.12}\pm 0.18$ 96.08 0.28 95.61 2.70
NAD[17] $91.30\pm 0.18$ $2.20\pm 0.43$ $88.88\pm 0.28$ $10.19\pm 12.98$ 94.25 2.12 95.24 29.42
NPD[37] $85.71\pm 3.59$ $3.39\pm 2.47$ $83.26\pm 3.59$ $5.26\pm 3.82$ 91.38 50.35 91.92 85.17
FT-SAM[36] $\textbf{92.23}\pm 0.12$ $2.37\pm 0.81$ $\textbf{90.33}\pm 0.15$ $8.12\pm 11.86$ 96.38 1.93 94.29 6.16
I-BAU[33] $88.78\pm 1.53$ $5.53\pm 6.81$ $86.16\pm 0.91$ $32.79\pm 18.81$ 25.96 10.87 12.91 68.68
ETF-FT(ours) $91.77\pm 0.16$ $2.85\pm 0.68$ $90.24\pm 0.28$ $4.22\pm 2.05$ 91.99 1.15 87.55 1.60
Table 2: Comparison of cleansing algorithms on trojaned models trained on 10% poisoned data with the BadNet attack. Results are reported on CIFAR-10. Results on ResNet18 are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
@@
Construct random ETF $\textbf{W}_{\text{ETF}}$
@@
PreAct-ResNet18[13] ViT-B/16[1]
5% Clean data 1% Clean data 5% 1%
Method ACC ASR ACC ASR ACC ASR ACC ASR
Original 91.17 99.90 91.17 99.90 96.67 100.00 96.67 100.00
FT $91.92\pm 0.15$ $3.12\pm 3.14$ $90.60\pm 0.37$ $3.75\pm 6.41$ 95.94 100.00 95.98 100.00
NCleanse[29] $91.49\pm 0.45$ $41.07\pm 48.04$ $90.50\pm 1.05$ $70.67\pm 44.64$ 96.80 2.33 96.52 40.96
NAD[17] $91.36\pm 0.28$ $1.51\pm 0.37$ $88.39\pm 0.89$ $2.99\pm 1.94$ 89.99 100.00 96.40 100.00
NPD[37] $82.65\pm 2.14$ $16.34\pm 11.34$ $86.87\pm 0.13$ $43.03\pm 1.68$ 96.69 100.00 96.50 100.00
FT-SAM[36] $90.78\pm 0.21$ $8.10\pm 4.62$ $87.52\pm 0.60$ $12.23\pm 19.80$ 96.98 100.00 96.82 100.00
I-BAU[33] $91.83\pm 0.12$ ${\color[rgb]{1,0,0}82.22}\pm 1.73$ $\textbf{91.73}\pm 0.25$ ${\color[rgb]{1,0,0}99.91}\pm 0.01$ 36.10 9.74 36.70 0.41
ETF-FT(ours) $\textbf{92.39}\pm 0.24$ $\textbf{0.04}\pm 0.05$ $91.01\pm 0.47$ $0\textbf{0.00}\pm 0$ 0.00 95.34 10.76 95.04 10.11
Table 3: Comparison of cleansing algorithms on trojaned models trained with 10% poisoned data with the TrojanNN attack. Results are reported on CIFAR-10. Results on ResNet are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
@@ 4.2 Experiment Details
We compare our method with a variety of state-of-the-art cleansing algorithms that have demonstrated strong performance on the BackdoorBench [30] leaderboard. As a baseline, we also run vanilla fine-tuning, with no trojan-aware modifications, on the same subset of clean data–these results are labeled FT in the following tables. Our evaluation considers two types of backdoor attacks: BadNets-A2O [10] and TrojanNN [19], and two model architectures: PreAct-ResNet18 [13] and ViT-B/16 [1].
We focus on two key metrics to evaluate the performance of each cleansing method: accuracy on clean data (ACC) and attack success sate (ASR). Accuracy reflects the model’s performance on clean data after cleansing, ideally showing no reduction compared to the original model. ASR, on the other hand, measures the proportion of poisoned samples classified as the attack target class. An effective cleansing method should reduce ASR to nearly zero.
We utilize the BackdoorBench repository [30] implementations of all attacks and cleansing methods besides our own. For experiments on PreAct-ResNet18, we use the original hyperparameters provided in the repository. By default, the repository uses 5% of the training data as available clean data for fine-tuning, though we also run experiments with 1% of training data and with imbalanced and corrupted data, as in real world situations the end user may not have access to (much of) the original training data.
@@ 4.3 Results
BadNet Attack Tab. 2 presents the cleansing results for the BadNet attack. When the cleansing algorithms are allowed to use 5% of the training data to fine-tune, on the ResNet model, ETF-FT achieves competitive ASR while maintaining high accuracy on clean data relative to the other methods tested. In more restrictive conditions, where only 1% of the training data is available, ETF-FT and Neural Cleanse maintain strong performance, while other methods show significant degradation. The numbers in red indicate extremely poor performance, failing to function as an effective cleansing method.
@@
Vision Transformers Despite their state-of-the-art performance and growing popularity, vision transformers have rarely been examined in the context of backdoor cleansing performance. This is in part due to the fact that their distinct architecture makes it difficult to adapt many common cleansing methods. Tab. 2 shows cleansing results for the BadNet attack applied to the ViT model, ETF-FT achieves very low ASR in both standard and limited data conditions, surpassing other methods except Neural Cleanse in the standard data case. This does come at the cost of a trade-off with in clean accuracy, which we believe can be mitigated, as discussed in Sec. 5.1.
@@
TrojanNN Attack Alongside the standard BadNet attack, we consider the more sophisticated TrojanNN attack [19] which identifies specific vulnerable neurons in a pre-trained network and reverse engineers a subtle trigger to target them. As shown in Tab. 3, ETF-FT is particularly well suited to cleansing this style of trojan. On ResNets, ETF-FT completely eliminates the trigger, achieving an ASR of nearly 0 across 10 trials, while maintaining competitive accuracy with the other cleansing methods. On ViT, many methods, implemented out of the box, fail completely to eliminate the trigger, while ETF-FT achieves an ASR of 10% with only a 1.5 point drop in clean accuracy. I-BAU algorithm is able to almost entirely eliminate the trigger, but in doing so destroyed the model’s accuracy on clean data.
@@ 4.3.1 Robustness
The above results show that ETF-FT performs well with access to only a small amount of training data. But it is likely that the user of a trojaned model may not have access to any data from the original training distribution at all. We simulated distribution shift in two ways: by imbalancing the class distribution and by applying random erasures within image patches. Results for these datasets, applied to the ResNet BadNet poisoned models, are shown in Tab. 4 and Tab. 8 (in Appendix A). In both settings, ETF-FT excels at maintaining the model’s clean accuracy along with competitive ASR relative to other methods. Neural Cleanse, which had generally very strong results in this setting with training data, performs significantly better than all other methods on the imbalanced data, but its clean accuracy suffers on the corrupted data. This is due to the fact that its pruning method relies on identifying important neurons based on their activations on the available data–so it is vulnerable to accidentally pruning neurons associated with valid features which aren’t represented in the fine-tuning dataset.
@@
NCleanse[29] $81.80\pm 0.38$ $1.72\pm 0.38$
NAD[17] $67.14\pm 1.96$ $3.10\pm 1.68$
NPD[37] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[36] $85.23\pm 0.77$ $\textbf{1.21}\pm 0.48$
I-BAU[33] $86.65\pm 1.38$ $28.26\pm 17.91$
@@
Table 4: Data Corruption: Comparison of cleansing algorithms with 1% corrupted (under random erasure) clean training data on BadNet trojaned models. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 5 random trials with mean±$\pm$±std.
@@ 4.3.2 More Sophisticated Trigger
To thoroughly evaluate the performance of ETF-FT against Trojan attacks with more sophisticated triggers beyond the basic yet powerful patch trigger, we conduct experiments using Instagram filter triggers [28]. A demonstration of the filter effects is presented in Fig. 7, where the Gotham filter introduces blurring and a gray tone, while the Lomo filter creates a vignette effect. As shown in Tab. 5, ETF-FT remains effective even against this more subtle type of trigger.
Figure 7: Instagram filter effect: Original (Top), Gotham (Middle), Lomo (Bottom)
Gotham Lomo
Method ACC ASR ACC ASR
@@
NCleanse[29] $84.54\pm 6.05$ ${\color[rgb]{1,0,0}70.49}\pm 28.77$ $83.1\pm 5.18$ $22.68\pm 38.73$
NAD[17] ${\color[rgb]{1,0,0}29.48}\pm 2.61$ $4.14\pm 4.03$ ${\color[rgb]{1,0,0}29.52}\pm 2.48$ $\textbf{0.0}\pm 0$ 0.0
NPD[37] $87.21\pm 2.53$ $\textbf{3.87}\pm 5.29$ $87.68\pm 3.62$ $20.82\pm 18.39$
FT-SAM[36] $85.94\pm 1.42$ $10.59\pm 12.85$ $86.93\pm 1.22$ $1.84\pm 1.03$
I-BAU[33] $\textbf{90.34}\pm 0.86$ $39.12\pm 27.04$ $89.86\pm 1.56$ $8.84\pm 5.58$
@@
Table 5: Performance comparison of cleansing algorithms on Trojaned models trained with 10% poisoned data using the BadNets attack, where the trigger is the Gotham and Lomo Instagram filters. Cleansing is performed with 5% clean data. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).

### v3

@@ 4.1 Methodology
This tension motivates an algorithm for cleansing a network of any potential trojan triggers, which we name ETF-FT. ETF-FT aims to take a fully trained, possibly trojaned neural network and remove trojan triggers, if any exist, while maintaining the network’s good performance on clean test data. The effects of a trojan attack are most apparent in fig. 3a, which shows that the weights of the final layer of a benign neural network exhibit a much greater degree of symmetry than those of an analogous trojaned network. So, as specified in algorithm 1, we proceed by first over-writing the weights of the final layer to a randomly generated simplex ETF, freezing those weights, and then fine-tuning the remaining model parameters on a small subset of clean data. Notably, this process works well with even only limited data, and the fine-tuning data does not necessarily need to come from the original training set.
@@
Construct random ETF $\textbf{W}_{\text{ETF}}$ according to algorithm 2
@@
Algorithm 2 Constructing the $\mathbf{W}_{\text{ETF}}$ Matrix
Input: Number of classes $K$, random orthogonal matrix $\mathbf{P}\in\mathbb{R}^{m\times K}$
Initialize: $\mathbf{I}\leftarrow\mathbf{I}_{K}$ (Identity matrix of size $K\times K$)
Initialize: $\mathbf{1}\leftarrow\mathbf{1}_{K\times K}$ (Matrix of ones of size $K\times K$)
Compute $\mathbf{W}_{\text{ETF}}\leftarrow\sqrt{\frac{K}{K-1}}\times\mathbf{P}\times\left(\mathbf{I}-\frac{1}{K}\mathbf{1}\right)^{T}$
Output: Matrix $\mathbf{W}_{\text{ETF}}$
@@ 4.2 Experiment Details
We focus on two key metrics to evaluate the performance of each cleansing method: accuracy on clean data (ACC) and attack success rate (ASR). Accuracy reflects the model’s performance on clean data after cleansing, ideally showing no reduction compared to the original model. ASR, on the other hand, measures the proportion of poisoned samples classified as the attack target class. An effective cleansing method should reduce ASR to nearly zero.
We compare our method against several state-of-the-art cleansing algorithms that have demonstrated strong performance on the BackdoorBench [wu2022backdoorbench] leaderboard. As a baseline we run vanilla fine-tuning, with no trojan-aware modifications, on the same subset of clean data–these results are labeled FT in our results. We utilize the BackdoorBench repository [wu2022backdoorbench] implementations of all attacks and cleansing methods besides our own. For experiments on PreAct-ResNet18, we adopt the default hyperparameters in the repository with a poisoning ratio 10% for all of our experiments. We evaluate a range of backdoor attacks under the BackdoorBench setting, including BadNets-A2O [badnets], TrojanNN [liu2018trojann], ReFool [liu2020reflection], WaNet [nguyen2021wanet], BPP [wang2022bppattack], CTRL [li2023embarrassingly], Blended [chen2017targeted], Blind [bagdasaryan2021blind], Input-aware [nguyen2020input], SIG [barni2019new], and LIRA [doan2021lira]. Among these, Blind, LIRA, and SIG fail to achieve an ASR above 60% while maintaining at least 70% ACC, and are thus excluded from further analysis. Notably, none of the existing cleansing methods effectively mitigate ReFool or Blended attacks—ASR remains above 40% when ACC is kept above 70%. Therefore, we primarily focus on the remaining attacks, especially BadNets-A2O and TrojanNN. Experiments are conducted on two model architectures: PreAct-ResNet18 [preactresnet] and ViT-B/16 [alexey2020image]. We attempted to evaluate on ImageNet to test scalability but were unable to train trojaned models with both high ASR ($\geq 50\%$) and reasonable ACC. Thus, these results are omitted.
(a) PreAct-ResNet18 Method 5% Benign 1% Benign ACC ASR ACC ASR Original 88.95 95.06 88.95 95.06 FT $91.63\pm 0.18$ $10.03\pm 7.23$ $90.20\pm 0.22$ $22.04\pm 25.38$ NCleanse[neuralCleanse] $91.57\pm 0.1$ $\textbf{0.98}\pm 0.12$ $89.90\pm 0.34$ $\textbf{1.12}\pm 0.18$ NAD[NeuralAttentionDistillation] $91.30\pm 0.18$ $2.20\pm 0.43$ $88.88\pm 0.28$ $10.19\pm 12.98$ NPD[zhu2024neural] $85.71\pm 3.59$ $3.39\pm 2.47$ $83.26\pm 3.59$ $5.26\pm 3.82$ FT-SAM[zhu2023enhancing] $\textbf{92.23}\pm 0.12$ $2.37\pm 0.81$ $\textbf{90.33}\pm 0.15$ $8.12\pm 11.86$ I-BAU[zeng2021adversarial] $88.78\pm 1.53$ $5.53\pm 6.81$ $86.16\pm 0.91$ $32.79\pm 18.81$ ETF-FT (ours) $91.77\pm 0.16$ $2.85\pm 0.68$ $90.24\pm 0.28$ $4.22\pm 2.05$
(b) ViT-B/16 Method 5% Benign 1% Benign ACC ASR ACC ASR Original 95.43 87.06 95.43 87.06 FT 96.06 20.46 95.74 36.87 NCleanse[neuralCleanse] 96.08 0.28 95.61 2.70 NAD[NeuralAttentionDistillation] 94.25 2.12 95.24 29.42 NPD[zhu2024neural] 91.38 50.35 91.92 85.17 FT-SAM[zhu2023enhancing] 96.38 1.93 94.29 6.16 I-BAU[zeng2021adversarial] 25.96 10.87 12.91 68.68 ETF-FT (ours) 91.99 1.15 87.55 1.60
Table 2: Comparison of cleansing methods against the BadNets-A2O attack on CIFAR-10 using 5% and 1% benign data. PreAct-ResNet18 results are averaged over 10 trials; ViT results are from a single trial.
@@ 4.3 Cleansing Results
BadNet Attack Table 2 presents the cleansing results for the BadNet attack. When the cleansing algorithms are allowed to use 5% of the training data to fine-tune, on the ResNet model, ETF-FT achieves competitive ASR while maintaining high accuracy on clean data relative to the other methods tested. In more restrictive conditions, where only 1% of the training data is available, ETF-FT and Neural Cleanse maintain strong performance, while other methods show significant degradation.
@@
Vision Transformers Despite their state-of-the-art performance and growing popularity, vision transformers have rarely been examined in the context of backdoor cleansing performance. This is in part due to the fact that their distinct architecture makes it difficult to adapt many common cleansing methods. Table 2 shows cleansing results for the BadNet attack applied to the ViT model, ETF-FT achieves very low ASR in both standard and limited data conditions, surpassing other methods except Neural Cleanse in the standard data case. This does come at the cost of a trade-off with in clean accuracy, which we believe can be mitigated, as discussed in section 5.1.
(a) PreAct-ResNet18
Method 5% Benign 1% Benign
ACC ASR ACC ASR
Original 91.17 99.90 91.17 99.90
FT $91.92\pm 0.15$ $3.12\pm 3.14$ $90.60\pm 0.37$ $3.75\pm 6.41$
NCleanse[neuralCleanse] $91.49\pm 0.45$ $41.07\pm 48.04$ $90.50\pm 1.05$ $70.67\pm 44.64$
NAD[NeuralAttentionDistillation] $91.36\pm 0.28$ $1.51\pm 0.37$ $88.39\pm 0.89$ $2.99\pm 1.94$
NPD[zhu2024neural] $82.65\pm 2.14$ $16.34\pm 11.34$ $86.87\pm 0.13$ $43.03\pm 1.68$
FT-SAM[zhu2023enhancing] $90.78\pm 0.21$ $8.10\pm 4.62$ $87.52\pm 0.60$ $12.23\pm 19.80$
I-BAU[zeng2021adversarial] $91.83\pm 0.12$ ${\color[rgb]{1,0,0}82.22}\pm 1.73$ $\textbf{91.73}\pm 0.25$ ${\color[rgb]{1,0,0}99.91}\pm 0.01$
ETF-FT (ours) $\textbf{92.39}\pm 0.24$ $\textbf{0.04}\pm 0.05$ $91.01\pm 0.47$ $\textbf{0.00}\pm 0.00$
(b) ViT-B/16
Method 5% Benign 1% Benign
ACC ASR ACC ASR
Original 96.67 100.00 96.67 100.00
FT 95.94 100.00 95.98 100.00
NCleanse[neuralCleanse] 96.80 2.33 96.52 40.96
NAD[NeuralAttentionDistillation] 89.99 100.00 96.40 100.00
NPD[zhu2024neural] 96.69 100.00 96.50 100.00
FT-SAM[zhu2023enhancing] 96.98 100.00 96.82 100.00
I-BAU[zeng2021adversarial] 36.10 9.74 36.70 0.41
ETF-FT (ours) 95.34 10.76 95.04 10.11
Table 3: Comparison of cleansing methods on TrojanNN-attacked CIFAR-10 models using 5% and 1% benign data. PreAct-ResNet18 results are averaged over 10 trials; ViT results are from a single trial.
@@
TrojanNN Attack Alongside the standard BadNet attack, we consider the more sophisticated TrojanNN attack [liu2018trojann] which identifies specific vulnerable neurons in a pre-trained network and reverse engineers a subtle trigger to target them. As shown in table 3, ETF-FT is particularly well suited to cleansing this style of trojan. On ResNets, ETF-FT completely eliminates the trigger, achieving an ASR of nearly 0 across 10 trials, while maintaining competitive accuracy with the other cleansing methods. On ViT, many methods, implemented out of the box, fail completely to eliminate the trigger, while ETF-FT achieves an ASR of 10% with only a 1.5 point drop in clean accuracy. I-BAU algorithm is able to almost entirely eliminate the trigger, but in doing so destroyed the model’s accuracy on clean data.
Other Attacks We also evaluate additional attacks, as detailed in section 4.2. Results for WaNet, BPP, and Input-aware are presented in table 4. ETF-FT demonstrates competitive performance against WaNet and BPP, and achieves state-of-the-art results on the Input-aware attack.
Method WaNet BPP Input-aware
ACC ASR ACC ASR ACC ASR
Original 93.69 97.12 94.58 99.98 94.25 99.36
FT 92.46 0.79 92.45 1.86 92.36 1.21
NCleanse[neuralCleanse] 92.85 39.19 87.51 21.57 86.07 12.79
NAD[NeuralAttentionDistillation] 91.68 0.88 31.55 8.95 32.63 5.85
NPD[zhu2024neural] 87.68 2.44 85.62 84.38 91.85 1.29
FT-SAM[zhu2023enhancing] 93.60 0.69 89.96 1.74 91.14 3.57
I-BAU[zeng2021adversarial] 91.85 4.24 91.84 5.21 92.08 10.35
ETF-FT (ours) 93.29 0.88 93.49 16.60 93.38 0.96
Table 4: Comparison of cleansing methods on WaNet, BPP, Input-aware-attacked CIFAR-10 models using 5% benign data. Results are on PreAct-ResNet18, averaged over 10 trials. Standard deviation in table 9 in appendix B.
@@ 4.3.1 Robustness
NCleanse[neuralCleanse] $81.80\pm 0.38$ $1.72\pm 0.38$
NAD[NeuralAttentionDistillation] $67.14\pm 1.96$ $3.10\pm 1.68$
NPD[zhu2024neural] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[zhu2023enhancing] $85.23\pm 0.77$ $\textbf{1.21}\pm 0.48$
I-BAU[zeng2021adversarial] $86.65\pm 1.38$ $28.26\pm 17.91$
@@
Table 5: Data Corruption: Comparison of cleansing algorithms with 1% corrupted (under random erasure) data on PreAct-ResNet18 trained on CIFAR-10 with a BadNet attack, averaged over 5 trials.
The above results show that ETF-FT performs well with access to only a small amount of training data, but it is likely that the user of a trojaned model may not have access to any data from the original training distribution at all. We simulated distribution shift in two ways: by imbalancing the class distribution and by applying random erasures within image patches. Results for these datasets, applied to the ResNet BadNet poisoned models, are shown in table 5 and table 8 (in appendix A). In both settings, ETF-FT excels at maintaining the model’s clean accuracy along with competitive ASR relative to other methods. Neural Cleanse, which had strong results with access to the original training data, performs significantly better than all other methods on the imbalanced data, but its clean accuracy suffers on the corrupted data. This is due to the fact that its pruning method relies on identifying important neurons based on their activations on the available data–so it is vulnerable to accidentally pruning neurons associated with valid features which aren’t represented in the fine-tuning dataset.
@@ 4.3.2 More Sophisticated Trigger
To thoroughly evaluate the performance of ETF-FT against Trojan attacks with more sophisticated triggers beyond the basic yet powerful patch trigger, we conduct experiments using Instagram filter triggers [trojai_instagram_xforms]. A demonstration of the filter effects is presented in section 4.3.2, where the Gotham filter introduces blurring and a gray tone, while the Lomo filter creates a vignette effect. As shown in table 6, ETF-FT remains effective even against this more subtle type of trigger.
[…]
figureInstagram filter effects (Original, Gotham, Lomo)
Method Gotham Lomo
ACC ASR ACC ASR
@@
NCleanse[neuralCleanse] $84.54\pm 6.05$ ${\color[rgb]{1,0,0}70.49}\pm 28.77$ $83.1\pm 5.18$ $22.68\pm 38.73$
NAD[NeuralAttentionDistillation] ${\color[rgb]{1,0,0}29.48}\pm 2.61$ $4.14\pm 4.03$ ${\color[rgb]{1,0,0}29.52}\pm 2.48$ $\textbf{0.0}\pm 0.0$
NPD[zhu2024neural] $87.21\pm 2.53$ $\textbf{3.87}\pm 5.29$ $87.68\pm 3.62$ $20.82\pm 18.39$
FT-SAM[zhu2023enhancing] $85.94\pm 1.42$ $10.59\pm 12.85$ $86.93\pm 1.22$ $1.84\pm 1.03$
I-BAU[zeng2021adversarial] $\textbf{90.34}\pm 0.86$ $39.12\pm 27.04$ $9.86\pm 1.56 8.8$$84\pm 5.58$
@@
Table 6: Performance of cleansing methods on BadNets-trojaned CIFAR-10 models using Gotham and Lomo Instagram filter triggers. Cleansing is performed with 5% clean data. Results are reported on PreAct-ResNet18, averaged over 10 trials.